Solve the initial value problem (2x – y2)dx + (1 – 2xy)dy = 0, y(1) = 5
Answer #1

0 (2x - y dx + (1-2xy) dy =0, y 12) = 5 Componing with Midat Nay = 2 We have ЭМ M = ax - y 행 N=1-2xy -2y ON 2x SMA ON 2x म 18